ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 154, 2006

Albano, P. and A. Bove Analytic stratifications and the cut locus of a class of distance functions

61

Bárány, I. and M. Prodromou On maximal convex lattice polygons inscribed in a plane convex set

337

Bartholdi, L. Branch rings, thinned rings, tree enveloping rings  93

Bonnet, R. and M. Rubin Chains of well-generated Boolean algebras whose union is not well-generated

141

Bove, A. See P. Albano

Dade, E. C. and M. K. Yadav Finite groups with many product conjugacy classes

29

Dekimpe, K. and H. Pouseele Cohomology of virtually nilpotent groups with coefficient in ℝ^{k}

1

Farah, I. Analytic Hausdorff gaps II: The density zero ideal  235

Izumiya, S., D. Pei and M. C. Romero Fuster The horospherical geometry of surfaces in hyperbolic 4-space

361

Janko, Z. Finite quaternion-free 2-groups  157
Janko, Z. Minimal non-quaternion-free finite 2-groups  185

Koldobsky, A., V. Yaskin and M. Yaskina Modified Busemann– Petty problem on sections of convex bodies

191

Lev, V. F. Large sum-free sets in ℤ/pℤ  221

van Mill, J. and R. Pol A complete C-space whose square is strongly infinite-dimensional

209

Moreno Mejía, I. The Chern classes of the eigenbundles of an automorphism of a curve

247

Mossel, E., R. O’Donnell, O. Regev, J. E. Steif and B. Sudakov Non-interactive correlation distillation, inhomogeneous Markov chains, and the reverse Bonami–Beckner inequality

299

O’Donnell, R. See E. Mossel

Offen, O. On symplectic periods of the discrete spectrum of GL_{2n}  253
Pan, H. and Z.-W. Sun Restricted sumsets and a conjecture of Lev  21

Pei, D. See S. Izumiya

Pol, R. See J. van Mill Pouseele, H. See K. Dekimpe Prodromou, M. See I. Bárány Regev, O. See E. Mossel Romero Fuster, M. C. See S. Izumiya Rubin, M. See R. Bonnet Steif, J. E. See E. Mossel Sudakov, B. See E. Mossel Sun, Z.-W. See H. Pan

Wiesend, G. Covers of varieties with completely split points  51

Yadav, M. K. See E. C. Dade Yaskin, V. See A. Koldobsky Yaskina, M. See A. Koldobsky